how many ml of that is 10% alcohol must be added to a 500 ml wine of 5% alcohol to make a wine that is 7.5% alcohol?
Answer by mananth(16946)

-------------- percent ---------------- Amount
Wine 5 ---------------- 500 ml
Alcohol 10 ---------------- x ml
Mixture 7.5 ---------------- 500 + x ml

5*500+10x=7.5(500+x)
2500+10x=3750+7.5x
10x-7.5x=-2500+3750
2.5x=1250
/ 2.5
x=500 ml Alcohol

m.ananth@hotmail.ca